OpenStudy (anonymous):

all these are linear equations are equivalent to 3x-4y=20 except for one:

OpenStudy (anonymous):

\[a. \frac{ 3 }{ 4 }x-5\]

OpenStudy (anonymous):

\[b. 8y+40=6x\]

OpenStudy (anonymous):

\[c.2y=\frac{ 3 }{ 2 }x+10\]

OpenStudy (anonymous):

C :P

OpenStudy (anonymous):

\[d. \frac{ 1 }{ 5 }y=\frac{ 3 }{ 20 }x-1\]

OpenStudy (anonymous):

@sami-21 can you kindly explain it to me? i have a big test tomorrow

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

simplify the give equation little bit \[\huge 3x-20=4y \] divide by 4 both sides \[\Huge y=\frac{3}{4}x-5\] ( this is in fact option a )

OpenStudy (anonymous):

now for option b divide the whole equation with 2 4y+20=3x or \[\Large 4y=3x-20\] divide both sides by 4 \[\Large y=\frac{3}{4}x-5\] which is again simila to a and given equation

OpenStudy (anonymous):

now check option d multiply whole equation with 5 \[\Large y=\frac{3}{4}x-5\] which is also similar to previous ones

OpenStudy (anonymous):

for option c divide whole equation with 2 \[\Large y=\frac{3}{4}x+5\] its different there is plus sign with 5 !!!

OpenStudy (anonymous):

Still confused ?

OpenStudy (anonymous):

a little bit.

OpenStudy (anonymous):

ok lets look at another way we will make the original equation from option a, b and d

OpenStudy (anonymous):

yayy!

OpenStudy (anonymous):

option a is \[\Large y=\frac{3}{4}x-5\] multiply whole equation with 4 \[\Large 4*y=4(\frac{3}{4})x-4*5\] \[\Large 4y=3x-20\] rearrange \[\Large 3x-4y =20\] are you ok with this ? let me know if there is any confusion

OpenStudy (anonymous):

i see so thats why you chose D because it has the only positive sign !

OpenStudy (anonymous):

i choosed C :)

OpenStudy (anonymous):

i mean c xD haha

OpenStudy (anonymous):

Sorry I got to go will be back after a while .

OpenStudy (anonymous):

okay sure ! thanks for the help btw

OpenStudy (anonymous):

welcome :)